# HAPO: Training Language Models to Reason Concisely via History-Aware Policy Optimization

### v1

@@ Abstract
While scaling the length of responses at test-time has been shown to markedly improve the reasoning abilities and performance of large language models (LLMs), it often results in verbose outputs and increases inference cost. Prior approaches for efficient test-time scaling, typically using universal budget constraints or query-level length optimization, do not leverage historical information from previous encounters with the same problem during training. We hypothesize that this limits their ability to progressively make solutions more concise over time. To address this, we present History-Aware Policy Optimization (HAPO), which keeps track of a history state (e.g., the minimum length over previously generated correct responses) for each problem. HAPO employs a novel length reward function based on this history state to incentivize the discovery of correct solutions that are more concise than those previously found. Crucially, this reward structure avoids overly penalizing shorter incorrect responses with the goal of facilitating exploration towards more efficient solutions. By combining this length reward with a correctness reward, HAPO jointly optimizes for correctness and efficiency. We use HAPO to train DeepSeek-R1-Distill-Qwen-1.5B, DeepScaleR-1.5B-Preview, and Qwen-2.5-1.5B-Instruct, and evaluate HAPO on several math benchmarks that span various difficulty levels. Experiment results demonstrate that HAPO effectively induces LLMs’ concise reasoning abilities, producing length reductions of 33-59% with accuracy drops of only 2-5%. [^1] Our code and data is available at https://github.com/HCY123902/HAPO.
**footnotetext: Equal contribution.
1 Introduction
@@
Recent advances in large language models (LLMs) have highlighted the effectiveness of test-time scaling [7, 15, 46, 22, 21, 36], which enables LLMs to develop longer and more sophisticated reasoning behaviors such as self-reflection and verification, substantially improving their performance across a wide range of tasks. While generating long reasoning chains can significantly boost models’ accuracy, it induces verbosity, increases inference cost, and incurs substantial computational and memory overhead due to the quadratic attention complexity and linear key-value (KV) cache growth that are inherent in transformer architectures [37]. Notably, even for very simple problems like “What is the answer to 2 plus 3?", these reasoning models tend to engage in unnecessarily lengthy reasoning and produce solutions that span hundreds or even thousands of tokens when only a few steps would suffice [17, 4, 34]. This inefficiency, often referred to as overthinking, limits the practical deployment of reasoning models in real-world applications.
@@
There have been initial explorations into limiting the generation length of reasoning models, either by enforcing a fixed, universal budget constraint to be applied to each question (i.e., universal budget forcing) [9, 23, 11, 1, 14], or by dynamically optimizing the reasoning length for each question [20, 2, 45, 28, 24] (i.e., query-level optimization).
@@
Universal budget forcing methods intervene when a predefined token limit is reached. For reinforcement learning (RL) based methods, this usually results in a non-positive reward. For solutions that use supervised-finetuning (SFT) or few-shot prompting, it implies forcing a response length cutoff by inserting an end-of-thinking token. For example, S1 [23] introduces a length control mechanism that injects special tokens (e.g., Final Answer) when responses are too long. Despite reducing output length with respect to a base model, such fixed token budgets cannot address the overthinking problem; instead they unnecessarily constrain exploratory reasoning on harder questions and still allow excessive token usage on simpler ones [9, 23, 11, 1, 14].
In contrast, query-level optimization methods learn a different budget for each query. For example, Arora and Zanette [2] compare the lengths of the responses to the same query in each training step and reward shorter responses while penalizing longer ones. Although these methods alleviate the overthinking problem to some extent, they cannot leverage historical information derived from previous encounters with the same query during training, i.e., across multiple training steps. This limits their ability to progressively make solutions more concise over time. For example, suppose responses sampled initially were considerably longer than those sampled in a subsequent encounter with the same problem. We argue that in this case, the model has become more efficient and, therefore, should receive a positive length reward in the later encounter.
@@
We use HAPO to train three open-source reasoning models (DeepSeek-R1-Distill-Qwen-1.5B [7], DeepScaleR-1.5B-Preview, and Qwen-2.5-1.5B-Instruct) and evaluate the models on math problems from three standard datasets—GSM8K [6], MATH500 [13], and AIME2024. Overall we find that HAPO effectively elicits LLMs’ concise reasoning abilities: when averaged across the three benchmarks, HAPO produces a 49% length reduction with only 2% accuracy drop for DeepSeek-R1-Distill-Qwen-1.5B; a 33% length reduction and 5% accuracy drop for DeepScaleR-1.5B-Preview; and a 59% length reduction and 2% accuracy drop for Qwen-2.5-1.5B-Instruct. Using DeepSeek-R1-Distill-Qwen-1.5B as the base model, we compare HAPO with baselines from universal budget forcing (the L1 models [1]) and query-level optimization (Query-Opt [2]) paradigms. Here, our results show that HAPO enables a better overall length-correctness trade-off. Notably, HAPO uses 19% fewer tokens than the most competitive baseline while achieving similar accuracy (see Figure 1).
2 Related Work
@@
Test-time scaling. Prior work [39, 32, 40] has shown that increasing test-time computation can improve the performance of large language models (LLMs), particularly in reasoning-intensive tasks such as math problem solving and code generation. Wei et al. [39] demonstrates that generating intermediate reasoning steps (i.e., chain-of-thought (CoT)) can effectively elicit the reasoning capabilities of LLMs, inspiring a wave of follow-up research. Many subsequent methods focus on parallel scaling, where multiple responses are sampled from the model and aggregated into a final answer, either through an external reward model or an internal voting mechanism. Representative approaches include Best-of-N sampling [40], beam search [32], majority voting [38] and its weighted variant [19], as well as tree-based search methods [44, 42, 3].
@@
Another line of work focuses on scaling the length of the response rather than the number of responses. It aims to enable LLMs to engage in more thorough reasoning and support self-correction and verification [18]. With the success of OpenAI’s o1 [15] model, more recent efforts have been made along this line. Notable efforts include DeepSeek’s R1 models [7], Qwen’s QwQ-series models, and other work that reproduces long CoT capabilities in smaller-scale models [46, 23]. A common feature of how most of these models are trained is the use of reinforcement learning (RL) algorithms [26, 27] and rule-based rewards. This training paradigm encourages models to generate increasingly long CoT in order to arrive at the correct answer.
@@
Efficient reasoning. While test-time scaling with long CoT significantly improves accuracy, it comes at the cost of computational inefficiency. In particular, reasoning models often produce verbose and unnecessary reasoning when solving simple problems—a phenomenon commonly referred to as overthinking [34].
@@
A variety of methods have been proposed to address this issue. For example, some approaches [10, 12, 5, 33, 29, 31] replace discrete vocabulary tokens with latent tokens in the continuous embedding space to perform reasoning. Other methods remain in the vocabulary space and focus on prompt optimization or additional training. Xu et al. [43] designs a prompt that explicitly encourages conciseness at each reasoning step, while Xia et al. [41], Han et al. [11], Shen et al. [30] employ offline training algorithms to teach LLMs to generate concise yet correct responses. However, such offline training typically requires carefully curated supervision data, which can be expensive and labor-intensive to obtain.
@@
In contrast, online RL training integrates length control directly within the reward signal. While the design of such length rewards varies across prior work, they generally fall into two categories: (1) Universal rewards, which compare response length against a predefined, universal budget and penalize responses exceeding this budget [14, 1], or utilize a universal penalty function [45]; (2) Query-level rewards, which are computed according to LLMs’ performance specific to each query. Many query-level rewards are comparison-based; some methods compare the lengths of responses to the same query within a batch, penalizing relatively longer ones [35, 2], while others compare the current response length against that of a reference model or ground truth [20, 28]. We argue that universal rewards are suboptimal due to their lack of adaptivity, necessitating a manual specification of budget estimates or penalty functions. Although query-level rewards seem more promising, current designs need improvement. For instance, in-batch comparison may not drive towards a global optimum (potentially rewarding long responses if all outputs in a batch are long), while comparison against a fixed reference can lead to rapid reward saturation once the model consistently outperforms the reference. In this work, we instead utilize the information from the training history to dynamically adjust the length reward per query, thereby providing the most up-to-date stimulus for the LLMs to compress their responses according to their own performance in the previous training steps.
3 History-Aware Length Reward
@@
In this section, we will first define our task setting (Section 3.1) and then describe our history-aware length reward (Section 3.2).
3.1 Task Setup
@@
Denote a $\theta$ parameterized LLM as $p_{\theta}$. Using the GRPO algorithm [^2] We explore other RL approaches such as PPO [26] in Appendix B. [27] we train $p_{\theta}$ on a dataset $\mathcal{D}=\{(x_{i},a^{*}_{i})\}_{i=1}^{N}$, where $x_{i}$ is a query (e.g., a math problem) and $a^{*}_{i}$ is its ground truth answer (e.g, a number). For each query $x_{i}$ in a batch, a set of responses $Y_{i}=\{y_{0},y_{1},...\}$ is sampled from the current LLM $p_{\theta}$. Following the procedure in [7], we can extract a candidate final answer $a_{i}$ from each response $y_{i}$. We train $p_{\theta}$ for multiple epochs; thus the model encounters each $x_{i}$ multiple times and we use $Y_{i}^{(j)}$ to refer to the set of responses generated by $p_{\theta}$ in its $j$-th encounter with $x_{i}$. Computation of the reward associated with $x_{i}$ is described next.
3.2 Reward Computation and History Update
@@
Accuracy reward. We employ a binary accuracy reward that compares the extracted answer $a_{i}$ against the ground truth $a^{*}_{i}$: $ra_{i}^{(j)}=1$ if $a_{i}^{(j)}=a_{i}^{*}$, and equals 00 otherwise.
@@
Length reward design. The length reward will be computed for each response $y_{i}^{(j)}\in Y_{i}^{(j)}$. At a high level, we desire a length reward with the following characteristics. (1) If $|y_{i}^{(j)}|<h_{i}$ and $a_{i}^{(j)}$ is correct, we should assign a positive reward, since $y_{i}^{(j)}$ is a more concise solution than any previous correct one. (2) If $|y_{i}^{(j)}|>h_{i}$ and $y_{i}^{(j)}$ is correct, then we should assign a negative reward. This is because $y_{i}^{(j)}$ is suboptimal as a shorter correct response has already been identified. (3) If $y_{i}^{(j)}$ is incorrect but $|y_{i}^{(j)}|<h_{i}$, we hypothesize that $p_{\theta}$ is exploring a potentially correct and shorter response and so propose to assign a neutral reward of 0. (4) If $|y_{i}^{(j)}|>h_{i}$ and $y_{i}^{(j)}$ is incorrect, then, similar to (2), we want a negative reward since $y_{i}^{(j)}$ is suboptimal. (5) In contrast to prior work [45, 30], we also want the reward to decrease smoothly as response length increases, even for incorrect responses, as long as $h_{i}$ is not Null. (6) The length reward should ensure that any correct response has an overall reward $r_{i}^{(j)}$ that is strictly higher than any incorrect one. (7) Lastly, if the reference $h_{i}$ is still Null, then this means $p_{\theta}$ has not generated any correct response in the training history, and therefore we give the current response a reward of 0 since no comparison can be made.
Figure 2 (right) depicts our length reward function that ranges from -1 to 1, rewards correct responses shorter than $h_{i}$, penalizes all responses longer than $h_{i}$, and smoothly decreases as $|y_{i}^{(j)}|$ increases. It relies on the cosine function in $f$ to meet characteristic (5) as well as a clipping cutoff $c\in[-1,0)$ for the $a_{i}^{(j)}=a^{*}_{i}$ case to meet characteristic (6):
@@
Final reward function. The final reward combines the length reward with the accuracy reward:
@@
$w$ is a hyperparameter within the range $[0,1]$ that controls the weight of the length component of the overall reward function.
History state update. After computing the reward for each response in $Y_{i}^{(j)}$, we update the history state $h_{i}$. Specifically, we extract the lengths of all correct responses in $Y_{i}^{(j)}$ as $L_{i}^{(j)}=\{|y_{i}^{(j)}||y_{i}^{(j)}\in Y_{i}^{(j)}\wedge a_{i}^{(j)}=a^{*}_{i}\}$. We then update $h_{i}$ using an aggregation function $aggre$ (i.e., $h_{i}\leftarrow aggre(\{h_{i}\}\cup L_{i}^{(j)})$). In this work, we set $aggre$ to the minimum function, so $h_{i}$ tracks the shortest correct response observed thus far. Other aggregation choices, such as taking the mean or median, are also possible.
4 Experiment Setup
@@
We train our models on 2,000 examples from the DeepScaleR-Preview-Dataset [22] and validate the models on another 500. The original dataset contains 40,000 math problems drawn from a variety of sources, spanning a difficulty spectrum from grade-school level (e.g., GSM8K [6]) to competitive mathematics (e.g., AMC, AIME). For the base models, we experiment with DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-R1-1.5B for short), DeepScaleR-1.5B-Preview (DeepScaleR-1.5B for short), and Qwen-2.5-1.5B-Instruct (Qwen-2.5-1.5B-Inst for short).
Both DeepSeek-R1-1.5B and DeepScaleR-1.5B are reasoning models, where the latter is an improved version of the former, with higher accuracy and shorter responses on reasoning tasks. In contrast, Qwen-2.5-1.5B-Inst is not a reasoning model, and is not trained specifically to generate long reasoning traces to solve complex questions. To apply HAPO to Qwen-2.5-1.5B-Inst, we curate an easier training set that comprises 2,000 examples from the train split of MATH dataset. This is to ensure that the model can still generate correct answers and receive positive length rewards for a reasonable number of queries during training. All models are trained for 5 epochs using the GRPO algorithm with HAPO hyperparameters set to $w=1.0$ and $c=-0.7$. More training details are in Appendix F.
@@
We evaluate the trained models on GSM8K [6], MATH500 [13], and AIME2024, which represent grade-school, intermediate, and competition-level math problems, respectively. Following DeepSeek-AI et al. [7], we use a sampling temperature of 0.6, top-$p$ of 0.95, and a maximum context length of 32,768 tokens. For GSM8K and MATH, we generate 4 responses per query, and for AIME2024, we generate 32 responses per query. As in prior work, we report Pass@1 (i.e., accuracy averaged across multiple sampled responses) and the average #Tokens in the responses for each benchmark.
@@
For universal budget forcing, we compare against L1-Exact and L1-Max [1]. Both methods employ GRPO with a token budget penalty alongside the accuracy reward. L1-Exact trains models to produce responses with exactly the target budget, while L1-Max treats the budget as an upper bound. To ensure a fair comparison—since the original models were trained on a significantly larger dataset—we rerun these methods using our own training data. For each benchmark, the L1-Exact budget is set to the average response length of HAPO, while L1-Max is given twice that average. Although exploring multiple token budgets for L1 baselines may improve their length-correctness trade-offs, we avoid this to maintain fairness and limit additional inference compute and human intervention.
For query-level optimization, we include the method proposed by Arora and Zanette [2], denoted as Query-Opt. This method compares the lengths of responses for the same query in each training batch, rewarding relatively shorter responses and penalizing longer ones.
5 Results and Analysis
We first present the main results (Section 5.1) on three models: DeepSeek-R1-1.5B, DeepScaleR-1.5B, and Qwen-2.5-1.5B-Inst. We then focus on DeepSeek-R1-1.5B for further experiments and analysis, including comparisons with prior work (Section 5.2), a case study on MATH500 (Section 5.3), out-of-domain evaluation (Section 5.4), and an inspection of training dynamics (Section 5.5). Finally, we examine the impact of various training hyperparameters, including the weight of the length reward (Section 5.6) and the size of the training set (Section 5.7).
5.1 Main Results
Table 1: Main results on GSM8K, MATH500, and AIME2024.
@@ 5.1 Main Results
Models Pass@1 #Tokens Pass@1 #Tokens Pass@1 #Tokens Pass@1 #Tokens
@@
On Qwen-2.5-1.5B-Instruct, HAPO again proves effective, reducing response length by 38%, 51%, and 62%, while keeping accuracy drops within 3%, 1%, and 1% on the three benchmarks. These results suggest that HAPO is broadly effective, even for models not originally optimized for reasoning.
Consistent with the findings in [2], we observe a reduction in token usage even when the model is trained solely with the correctness reward (i.e., HAPO w/ $w=0$). We hypothesize that this may be due to the relatively low difficulty of questions in the training set, though further investigation is needed to confirm the underlying cause. Compared to this baseline, HAPO produces responses that are, on average, 31% shorter and 4% less accurate on DeepSeek-R1-1.5B across all benchmarks. The corresponding reductions for DeepScaleR-1.5B and Qwen-2.5-1.5B-Inst are (26%, 5%) and (18%, 2%), respectively. These results suggest that HAPO consistently achieves a more favorable length-accuracy trade-off than optimizing for correctness alone.
5.2 Comparison with Prior Work
@@
Models Pass@1 #Tokens Pass@1 #Tokens Pass@1 #Tokens Pass@1 #Tokens

### v2

@@ Abstract
While scaling the length of responses at test-time has been shown to markedly improve the reasoning abilities and performance of large language models (LLMs), it often results in verbose outputs and increases inference cost. Prior approaches for efficient test-time scaling, typically using universal budget constraints or query-level length optimization, do not leverage historical information from previous encounters with the same problem during training. We hypothesize that this limits their ability to progressively make solutions more concise over time. To address this, we present History-Aware Policy Optimization (HAPO), which keeps track of a history state (e.g., the minimum length over previously generated correct responses) for each problem. HAPO employs a novel length reward function based on this history state to incentivize the discovery of correct solutions that are more concise than those previously found. Crucially, this reward structure avoids overly penalizing shorter incorrect responses with the goal of facilitating exploration towards more efficient solutions. By combining this length reward with a correctness reward, HAPO jointly optimizes for correctness and efficiency. We use HAPO to train DeepSeek-R1-Distill-Qwen-1.5B, DeepScaleR-1.5B-Preview, and Qwen-2.5-1.5B-Instruct, and evaluate HAPO on several math benchmarks that span various difficulty levels. Experiment results demonstrate that HAPO effectively induces LLMs’ concise reasoning abilities, producing length reductions of 33-59% with accuracy drops of only 2-5%.
Codehttps://github.com/HCY123902/HAPO
Introduction
@@
Recent advances in large language models (LLMs) have highlighted the effectiveness of test-time scaling (deepseek2025; open2024; zeng2025; deepscaler2025; deepcoder2025; qwq32b), which enables LLMs to develop longer and more sophisticated reasoning behaviors such as self-reflection and verification, substantially improving their performance across a wide range of tasks. While generating long reasoning chains can significantly boost models’ accuracy, it induces verbosity, increases inference cost, and incurs substantial computational and memory overhead due to the quadratic attention complexity and linear key-value (KV) cache growth that are inherent in transformer architectures (vaswani2017). Notably, even for very simple problems like “What is the answer to 2 plus 3?”, these reasoning models tend to engage in unnecessarily lengthy reasoning and produce solutions that span hundreds of tokens when only a few steps would suffice (kumar2025overthink; chen2024not; sui2025). This inefficiency, often referred to as overthinking, limits the practical deployment of reasoning models in real-world applications.
@@
There have been initial explorations into limiting the generation length of reasoning models, either by enforcing a fixed, universal budget constraint to be applied to each question (i.e., universal budget forcing) (fu2024efficiently; muennighoff2024; han2024; aggarwal2025; hou2025), or by dynamically optimizing the reasoning length for each question (luo2025; arora2025; yeo2025; she2025; qu2025) (i.e., query-level optimization).
@@
Universal budget forcing methods intervene when a predefined token limit is reached. For reinforcement learning (RL) based methods, this usually results in a non-positive reward. For solutions that use supervised-finetuning (SFT) or few-shot prompting, it implies forcing a response length cutoff by inserting an end-of-thinking token. For example, S1 (muennighoff2024) introduces a length control mechanism that injects special tokens (e.g., Final Answer) when responses are too long. Despite reducing output length with respect to a base model, such fixed token budgets cannot address the overthinking problem; instead they unnecessarily constrain exploratory reasoning on harder questions and still allow excessive token usage on simpler ones.
In contrast, query-level optimization methods learn a different budget for each query. For example, arora2025 compare the lengths of the responses to the same query in each training step and reward shorter responses while penalizing longer ones. Although these methods alleviate the overthinking problem to some extent, they cannot leverage historical information derived from previous encounters with the same query during training, i.e., across multiple training steps. This limits their ability to progressively make solutions more concise over time. For example, suppose responses sampled initially were considerably longer than those sampled in a subsequent encounter with the same problem. We argue that in this case, the model has become more efficient and, therefore, should receive a positive length reward in the later encounter. Unfortunately, existing approaches—including both universal budget forcing and query-level optimization methods—lack the mechanisms to utilize such historical signals.
@@
We use HAPO to train three open-source reasoning models, including DeepSeek-R1-Distill-Qwen-1.5B (deepseek2025), DeepScaleR-1.5B-Preview, and Qwen-2.5-1.5B-Instruct, and evaluate the models on math problems from three standard datasets—GSM8K (cobbe2021), MATH500 (hendrycks2021), and AIME2024. Overall we find that HAPO effectively elicits LLMs’ concise reasoning abilities: when averaged across the three benchmarks, HAPO produces a 49% length reduction with only 2% accuracy drop for DeepSeek-R1-Distill-Qwen-1.5B; a 33% length reduction with 5% accuracy drop for DeepScaleR-1.5B-Preview; and a 59% length reduction with 2% accuracy drop for Qwen-2.5-1.5B-Instruct. Using DeepSeek-R1-Distill-Qwen-1.5B as the base model, we compare HAPO with baselines from universal budget forcing (the L1 models (aggarwal2025)) and query-level optimization (Query-Opt (arora2025)) paradigms. Here, our results show that HAPO enables a better overall length-correctness trade-off. Notably, HAPO uses 19% fewer tokens than the most competitive baseline while achieving similar accuracy (see Figure 1).
Related Work
Test-time scaling
@@
Prior work (wei2022; snell2025; wu2024) has shown that increasing test-time computation can improve the performance of large language models (LLMs), particularly in reasoning-intensive tasks such as math problem solving and code generation. wei2022 demonstrates that generating intermediate reasoning steps (i.e., chain-of-thought (CoT)) can effectively elicit the reasoning capabilities of LLMs, inspiring a wave of follow-up research. Many subsequent methods focus on parallel scaling, where multiple responses are sampled from the model and aggregated into a final answer, either through an external reward model or an internal voting mechanism. Representative approaches include Best-of-N sampling (wu2024), beam search (snell2025), majority voting (wang2023) and its weighted variant (li2023), as well as tree-based search methods (yao2023; xin2025; besta2024).
@@
Another line of work focuses on scaling the length of the response rather than the number of responses. It aims to enable LLMs to engage in more thorough reasoning and support self-correction and verification (kumar2025). With the success of OpenAI’s o1 (open2024) model, more recent efforts have been made along this line. Notable efforts include DeepSeek’s R1 models (deepseek2025), Qwen’s QwQ-series models, and other work that reproduces long CoT capabilities in smaller-scale models (zeng2025; muennighoff2024). A common feature of how most of these models are trained is the use of reinforcement learning (RL) algorithms (schulam2017ppo; shao2024) and rule-based rewards. This training paradigm encourages models to generate increasingly long CoT in order to arrive at the correct answer.
Efficient reasoning
@@
While test-time scaling with long CoT significantly improves accuracy, it comes at the cost of computational inefficiency. In particular, reasoning models often produce verbose and unnecessary reasoning when solving simple problems—a phenomenon commonly referred to as overthinking (sui2025). Empirically, we observe that reasoning models such as DeepSeek-R1-Distill-Qwen-1.5B can generate over 6K tokens when solving problems from the MATH dataset (hendrycks2021), whereas human-written solutions only have around 200 tokens on average.
@@
A variety of methods have been proposed to address this issue. For example, some approaches (geiping2025; hao2024; cheng2024; su2025; shen2025efficient; shen2025codi) replace discrete vocabulary tokens with latent tokens in the continuous embedding space to perform reasoning. Other methods remain in the vocabulary space and focus on prompt optimization or additional training. xu2025 designs a prompt that explicitly encourages conciseness at each reasoning step, while xia2025; han2024; shen2025dast employ offline training algorithms to teach LLMs to generate concise yet correct responses. However, such offline training typically requires carefully curated supervision data, which can be expensive and labor-intensive to obtain.
@@
In contrast, online RL training integrates length control directly within the reward signal. While the design of such length rewards varies across prior work, they generally fall into two categories: (1) Universal rewards, which compare response length against a predefined, universal budget and penalize responses exceeding this budget (hou2025; aggarwal2025), or utilize a universal penalty function (yeo2025); (2) Query-level rewards, which are computed according to LLMs’ performance specific to each query. Many query-level rewards are comparison-based; some methods compare the lengths of responses to the same query within a batch, penalizing relatively longer ones (kimi2025; arora2025), while others compare the current response length against that of a reference model or ground truth (luo2025; she2025). We argue that universal rewards are suboptimal due to their lack of adaptivity, necessitating a manual specification of budget estimates or penalty functions. Although query-level rewards seem more promising, current designs need improvement. For instance, in-batch comparison may not achieve a global optimum (potentially rewarding long responses if all outputs in a batch are long), while comparison against a fixed reference can lead to rapid reward saturation once the model consistently outperforms the reference.
In this work, we instead follow the idea of utilizing past information from training history (zhang2023; shinn2023; le2025), to dynamically adjust the length reward per query, thereby providing the most up-to-date stimulus for the LLMs to compress their responses according to their own performance in the previous training steps.
History-Aware Length Reward
@@
In this section, we will first define our task setting and then describe our history-aware length reward.
Task Setup
@@
Denote a $\theta$ parameterized LLM as $p_{\theta}$. Using the GRPO algorithm [^1] We explore other RL approaches such as PPO (schulam2017ppo) in the Appendix. (shao2024) we train $p_{\theta}$ on a dataset $\mathcal{D}=\{(x_{i},a^{*}_{i})\}_{i=1}^{N}$, where $x_{i}$ is a query (e.g., a math problem) and $a^{*}_{i}$ is its ground truth answer (e.g, a number). For each query $x_{i}$ in a batch, a set of responses $Y_{i}=\{y_{0},y_{1},$...\} is sampled from the current LLM $p_{\theta}$. Following the procedure in (deepseek2025), we extract a candidate final answer $a_{i}$ from each response $y_{i}$. We train $p_{\theta}$ for multiple epochs; thus the model encounters each $x_{i}$ multiple times and we use $Y_{i}^{(j)}$ to refer to the set of responses generated by $p_{\theta}$ in its $j$-th encounter with $x_{i}$. Computation of the reward associated with $x_{i}$ is described next.
Reward Computation and History Update
@@
Accuracy reward
We employ a binary accuracy reward that compares the extracted answer $a_{i}$ against the ground truth $a^{*}_{i}$: $ra_{i}^{(j)}=1$ if $a_{i}^{(j)}=a_{i}^{*}$, and equals 0 otherwise.
@@
Length reward design
The length reward will be computed for each response $y_{i}^{(j)}\in Y_{i}^{(j)}$. At a high level, we desire a length reward with the following characteristics. (1) If $|y_{i}^{(j)}|<h_{i}$ and $a_{i}^{(j)}$ is correct, we should assign a positive reward, since $y_{i}^{(j)}$ is a more concise solution than any previous correct one. (2) If $|y_{i}^{(j)}|>h_{i}$ and $y_{i}^{(j)}$ is correct, then we should assign a negative reward. This is because $y_{i}^{(j)}$ is suboptimal with the presence of a shorter correct response. (3) If $y_{i}^{(j)}$ is incorrect but $|y_{i}^{(j)}|<h_{i}$, we hypothesize that $p_{\theta}$ is exploring a potentially correct and shorter response and so assign a neutral reward of 0. (4) If $|y_{i}^{(j)}|>h_{i}$ and $y_{i}^{(j)}$ is incorrect, then, similar to (2), we want a negative reward since $y_{i}^{(j)}$ is suboptimal. (5) In contrast to prior work (yeo2025; shen2025dast), we also want the reward to decrease smoothly as response length increases, even for incorrect responses, as long as $h_{i}$ is not Null. (6) The length reward should ensure that any correct response has an overall reward $r_{i}^{(j)}$ that is strictly higher than any incorrect one. (7) Lastly, if the reference $h_{i}$ is still Null, then this means $p_{\theta}$ has not generated any correct response in the training history, and therefore we give the current response a reward of 0 since no comparison can be made.
Figure 2 (right) depicts our length reward function that ranges from -1 to 1, rewards correct responses shorter than $h_{i}$, penalizes all responses longer than $h_{i}$, and smoothly decreases as $|y_{i}^{(j)}|$ increases. It relies on the cosine function in $f$ to meet characteristic (5) as well as a clipping cutoff $c\in(-1,0)$ for the $a_{i}^{(j)}=a^{*}_{i}$ case to meet characteristic (6):
@@
Final reward function
The final reward combines the length reward with the accuracy reward:
@@
$w$ is a hyperparameter within the range $[0,1]$ that controls the weight of the length component of the overall reward.
History state update
After computing the reward for each response in $Y_{i}^{(j)}$, we update the history state $h_{i}$. Specifically, we extract the lengths of all correct responses in $Y_{i}^{(j)}$ as $L_{i}^{(j)}=\{|y_{i}^{(j)}||y_{i}^{(j)}\in Y_{i}^{(j)}\wedge a_{i}^{(j)}=a^{*}_{i}\}$. We then update $h_{i}$ using an aggregation function aggreaggre (i.e., $h_{i}\leftarrow aggre(\{h_{i}\}\cup L_{i}^{(j)})$). In this work, we set aggreaggre to the minimum function, so $h_{i}$ tracks the shortest correct response observed thus far. Other aggregation choices, such as taking the mean or median, are also possible.
Experiment Setup
@@
We train our models on 2,000 examples from the DeepScaleR-Preview-Dataset (deepscaler2025) and validate the models on another 500. The original dataset contains 40,000 math problems drawn from a variety of sources, spanning a difficulty spectrum from grade-school level (e.g., GSM8K (cobbe2021)) to competitive mathematics (e.g., AMC, AIME). For the base models, we experiment with DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-R1-1.5B for short), DeepScaleR-1.5B-Preview (DeepScaleR-1.5B for short), and Qwen-2.5-1.5B-Instruct (Qwen-2.5-1.5B-Inst for short).
Both DeepSeek-R1-1.5B and DeepScaleR-1.5B are reasoning models, where the latter is an improved version of the former, with higher accuracy and shorter responses on reasoning tasks. In contrast, Qwen-2.5-1.5B-Inst is not a reasoning model, and is not trained specifically to generate long reasoning traces to solve complex questions. To apply HAPO to Qwen-2.5-1.5B-Inst, we curate an easier training set that comprises 2,000 examples from the train split of MATH dataset. This is to ensure that the model can still generate correct answers and receive positive length rewards for a reasonable number of queries during training. All models are trained for 5 epochs using the GRPO algorithm with HAPO hyperparameters set to $w=1.0$ and c=0.7c=-0.7. More training details are in the Appendix.
@@
We evaluate the trained models on GSM8K (cobbe2021), MATH500 (hendrycks2021), and AIME2024, which represent grade-school, intermediate, and competition-level math problems, respectively. Following deepseek2025, we use a sampling temperature of 0.6, top-$p$ of 0.95, and a maximum context length of 32,768 tokens. For GSM8K and MATH, we generate 4 responses per query, and for AIME2024, we generate 32 responses per query. As in prior work, we report Pass@1 (i.e., accuracy averaged across multiple sampled responses) and the average #Tokens in the responses for each benchmark.
@@
For universal budget forcing, we compare against L1-Exact and L1-Max (aggarwal2025). Both methods employ GRPO with a token budget penalty alongside the accuracy reward. L1-Exact trains models to produce responses with exactly the target budget, while L1-Max treats the budget as an upper bound. To ensure a fair comparison—since the original models were trained on a significantly larger dataset—we rerun these methods using our own training data. For each benchmark, the L1-Exact budget is set to the average response length of HAPO, while L1-Max is given twice that average. Although exploring multiple token budgets for L1 baselines may improve their length-correctness trade-offs, we avoid this to limit compute and human interventions, and maintain fairness.
For query-level optimization, we include the method proposed by arora2025, denoted as Query-Opt. This method compares the lengths of responses for the same query in each training batch, rewarding relatively shorter responses and penalizing longer ones. Additionally, we evaluate O1-Pruner (luo2025), which uses the base model’s responses to each query as a reference, to compute length and correctness rewards.
Results and Analysis
We first present the main results on three models: DeepSeek-R1-1.5B, DeepScaleR-1.5B, and Qwen-2.5-1.5B-Inst. We then focus on DeepSeek-R1-1.5B for further experiments and analysis, including comparisons with prior work, a case study on MATH500, and an inspection of training dynamics. Finally, we examine the impact of training set size on the model’s performance.
Main Results
@@
Models Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$
@@
On Qwen-2.5-1.5B-Inst, HAPO again proves effective, reducing response length by 38%, 51%, and 62%, while keeping accuracy drops within 3%, 1%, and 1% on the three benchmarks. These suggest that HAPO is broadly effective, even for models not originally optimized for reasoning.
Consistent with the findings in (arora2025), we observe a reduction in token usage even when the model is trained solely with the correctness reward (i.e., HAPO w/ $w=0$). We hypothesize that this may be due to the relatively low difficulty of questions in the training set, though further investigation is needed to confirm the underlying cause. Compared to this baseline, HAPO produces responses that are, on average, 31% shorter and 4% less accurate on DeepSeek-R1-1.5B across all benchmarks. The corresponding reductions for DeepScaleR-1.5B and Qwen-2.5-1.5B-Inst are (26%, 5%) and (18%, 2%), respectively. These suggest that HAPO achieves a more favorable length-accuracy trade-off than optimizing for correctness alone.
Comparison with Prior Work
@@
Models Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$